# On Secret-Message Transmission by Echoing Encrypted Probes

### v1

@@ VI Additional Comments
An important connection between STEEP and the works in [7], [2], [28] and [22] has been stressed in the introduction. The idea of feedback for WTC shown in [27] however has a limited applicability unlike STEEP.
@@
To achieve the secrecy capacity of STEEP, WTC coding is needed in the echoing phase (phase 2). Such coding schemes are available, including LDPC codes [19] and polar codes [18]. However, the current practicality of those codes due to their complexity seems questionable.
@@
STEEP may also remind one of a widely used method for networking security called “nonce”. The usefulness of nonce is based on the assumption that Alice can send a nonce reliably to Bob while Eve can not receive it. Then this nonce can be used (normally once) by Bob to encrypt a message to be sent to Alice. Unlike nonce, STEEP allows Eve to receive the probes from Alice but with some noise while Bob does not have to receive the probes with more accuracy than Eve, and the noisy probes received by Bob are used to encrypt a secret message to be sent to Alice. STEEP is naturally applicable at the physical layer due to presence of independent noises while it is also applicable at a higher layer.
@@ VII Conclusion
This paper has presented STEEP in its latest forms, and demonstrated a broad applicability of STEEP for secure communications especially at the physical layer. Its connection to some prior works such as in [7], [2], [28], [22] has been highlighted. Novel insights into how STEEP can be further formulated in different settings have been revealed. Furthermore, the achievable secrecy rates of STEEP in three different settings of practical importance have been derived and analyzed. In each of the settings, STEEP is shown to yield a positive secrecy rate, subject to a sufficiently large power in the echoing phase, as long as Eve’s receive channel in the probing phase is not noiseless. This attractive property of STEEP has its roots in the encryption lemma discussed in the introduction. Furthermore, the broad applicability of STEEP should make it a strong candidate for secure communications in practice - towards a full realization of a hope in [7].
@@
We will assume $n_{A}\geq n_{B}$. Recall $\mathbf{R}_{\Delta\mathbf{p}^{\prime}}$ in (III-B), $\mathbf{R}_{\Delta\mathbf{p}}$ in (III-A) and $C_{A|B,G}=\log\frac{N_{A|B}}{D_{A|B}}$ in (III-C). Also recall $\mathbf{H}_{AB}^{\prime\prime}=\sqrt{p_{B}/(2n_{B})}\mathbf{H}_{AB}$, $\mathbf{H}_{EB}^{\prime\prime}=\sqrt{p_{B}/(2n_{B})}\mathbf{H}_{EB}$, $\mathbf{H}_{BA}^{\prime}=\sqrt{p_{A}/n_{A}}\mathbf{H}_{BA}$, and $\mathbf{H}_{EA}^{\prime}=\sqrt{p_{A}/n_{A}}\mathbf{H}_{EA}$.
@@
$\lim_{p_{A}\to\infty}\log D_{A|B}$
$=\log\left|\frac{\eta_{p}n_{A}}{2n_{B}}\mathbf{H}_{AB}^{H}\mathbf{H}_{AB}(\boldsymbol{\Pi}_{BA}^{2})^{-1}+\mathbf{I}_{n_{B}}\right|$ (112)
@@
$\lim_{p_{A}\to\infty}\log N_{A|B}$
$=\lim_{p_{A}\to\infty}\log\left|\frac{p_{B}}{2n_{B}}\mathbf{H}_{AB}^{H}\mathbf{H}_{AB}\right|+o(\log p_{A}).$ (113)
@@
Now let us consider $C_{E|B,G}=\log\frac{N_{E|B}}{D_{E|B}}$ in (III-D) and $\mathbf{T}$ in (III-D). We know that $rank(\mathbf{H}_{EA})=\min(n_{A},n_{E})\doteq r_{A}$. We can write the eigenvalue decomposition (EVD) of $\mathbf{H}_{EA}^{H}\mathbf{H}_{EA}$ as $\mathbf{U}\mathbf{D}^{2}\mathbf{U}^{H}=\mathbf{U}_{A}\mathbf{D}_{A}^{2}\mathbf{U}_{A}^{H}$ where $\mathbf{D}_{A}^{2}$ is $r_{A}\times r_{A}$ nonsingular diagonal and $\mathbf{U}_{A}$ is the corresponding $r_{A}$ columns of the $n_{A}\times n_{A}$ unitary matrix $\mathbf{U}$. It follows from (III-D) that
@@
$C_{key}=\log\left|\mathbf{I}_{n_{A}}+\frac{p_{A}}{n_{A}}\mathbf{\tilde{H}}_{EA}^{H}\mathbf{\tilde{H}}_{EA}\right|$
$\,\,-\log\left|\frac{p_{A}}{n_{A}}\mathbf{H}_{EA}^{H}\mathbf{H}_{EA}+\mathbf{I}_{n_{A}}\right|$ (128)
@@
Next we consider $C_{s,G}^{B,A}\doteq\lim_{p_{A}\to\infty}C_{s,G}^{B}$. Since $\mathbf{H}_{EA}^{H}\mathbf{H}_{EA}$ is invertible, then $\lim_{p_{A}\to\infty}\mathbf{T}=\mathbf{I}_{n_{B}}$ and $\lim_{p_{A}\to\infty}\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}=0$. Also note that $\mathbf{R}_{\mathbf{\hat{p}}}=\mathbf{I}_{n_{B}}+\mathcal{O}(1/p_{A})$ and $\lim_{p_{A}\to\infty}\mathbf{R}_{\Delta\mathbf{p}^{\prime}}=\lim_{p_{A}\to\infty}\mathbf{R}_{\Delta\mathbf{p}}=0$. Therefore,
$C_{s,G}^{B,A}\doteq\lim_{p_{A}\to\infty}C_{s,G}^{B}$
$=\lim_{p_{A}\to\infty}\log\left(\frac{|\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}|}{|\mathbf{R}_{\Delta\mathbf{p}}|}\right)$ (132)
@@
$C_{s,G}^{B,A}=\lim_{p_{A}\to\infty}\log|\mathbf{R}_{\Delta\mathbf{p}}^{-1}\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}|$
@@
It is easy to verify that subject to (107), $C_{s,1}>0$ if and only if
@@
We will also show explicitly that $t_{1,M}<M-1$. This means that if $S_{E,A}=\alpha_{1}S_{1}>M-2$, then $01-\frac{t_{1,M}}{\alpha_{1}S_{1}+1}>$ 0 and hence (138) is equivalent to (108). Since $C_{s,1}$ must be a decreasing function of $\alpha_{1}$ (which is the ratio of the channel strength from AP to Eve over that from AP to UE1), $C_{s,1}$ must increase as $\alpha_{1}$ decreases. If the peak value of $\frac{t_{1,M}}{\alpha_{1}S_{1}+1}$ (a decreasing function of $\alpha_{1}$) is larger than one, then as $\alpha_{1}$ decreases (starting from the condition $S_{E,A}=\alpha_{1}S_{1}>M-2$) the condition $01-\frac{t_{1,M}}{\alpha_{1}S_{1}+1}>$ 0 would be reversed and a contradiction from $C_{s,1}>0$ would be concluded. Therefore, $t_{1,M}<\alpha_{1}S_{1}+1$ must hold for all $\alpha_{1}$. This also means that $t_{1,M}<1$ when $S_{E,A}=\alpha_{1}S_{1}=0$.
@@
Also recall $c_{i}=\sigma_{\hat{p}_{i}}^{2}=\mathcal{E}\{|\hat{p}_{i}|^{2}\}$; and for $n_{A}=1$, $\phi_{i,j}=1$ and hence $\epsilon_{i,j}=c_{i}c_{j}$.
@@
$q_{b}=-\mathbf{c}_{a}^{H}\mathbf{B}_{1,1}\mathbf{c}_{a}\mathbf{g}_{A}^{H}(\mathbf{g}_{A}\mathbf{g}_{A}^{H}+\mathbf{I})^{-1}\mathbf{g}_{A}$
$=-q_{a}\frac{S_{E,A}}{S_{E,A}+1}.$ (159)
@@
$\gamma_{1}-1=c_{1}-q_{1}$
$=\left(c_{1}-c_{1}^{2}\frac{S_{E,A}}{S_{E,A}+1}\right)-\frac{c_{1}^{2}t_{1,M}}{(S_{E,A}+1)^{2}}$ (184)

### v2

@@ VI Additional Comments
An important connection between STEEP and the works in [7], [2], [28] and [22] has been stressed in the introduction. The idea of feedback for WTC shown in [27] however has a limited applicability unlike STEEP. The secure feedback channel required in [36] and many other prior works is not required by STEEP.
@@
To achieve the achievable secrecy rate of STEEP, WTC coding is needed in the echoing phase (phase 2). Such coding schemes are available, including LDPC codes [19] and polar codes [18]. However, the current practicality of those codes due to their complexity seems questionable.
@@
STEEP may also remind one of a widely used method for networking security called “nonce”. The usefulness of nonce is based on the assumption that Alice can send a nonce reliably to Bob while Eve can not receive it. Then this nonce can be used (normally once) by Bob to encrypt a message to be sent to Alice. Unlike nonce, STEEP allows Eve to receive the probes from Alice but with some noise while Bob does not have to receive the probes with more accuracy than Eve, and the noisy probes received by Bob are used to encrypt a secret message to be sent to Alice. STEEP is naturally applicable at the physical layer due to presence of independent noises (especially thermal noises) while its applicability at a higher layer is also of great interest.
@@ VII Conclusion
Although related to some contributions in [7], [2], [28] and [22], STEEP as shown in this paper has a broad applicability for secure communications at the physical layer. This paper has presented: STEEP based on Gaussian probing signals and Gaussian linear encryption over Gaussian MIMO channels (G-STEEP); STEEP based on phase-shift-keying (PSK) probing signals and nonlinear PSK encryption over Gaussian SISO channel (P-STEEP); and a special form of G-STEEP for orthogonal multiple access between an access point and multiple users (M-STEEP). Achievable secrecy rates of these schemes have been derived and analyzed. It has been shown that positive secrecy rates for both single-link problem and multiple-access problem can be virtually guaranteed by asymmetric power allocation as long as Eve’s receive channel in the probing phase of STEEP is not noiseless, which includes the secrecy rate from a user to AP subject to exposure of messages from all other users. Such a discovery is highly novel, and in great contrast to numerous works in the physical layer security literature over past three decades that require user’s (including AP here and below) receive channel being stronger than Eve’s, user’s antennas more than Eve’s, reciprocal channel responses between users, secure feedback channel between users, collaborative third party (such as relay), and/or in-band full-duplex between users, in order to ensure a positive secrecy rate between users. While rooted in the encryption lemma discussed in the introduction, STEEP that exploits echoing encrypted probes, asymmetric power allocation and/or collaborative round-trip coding should have opened a new door of research and development for secure communications.
@@
We will assume $n_{A}\geq n_{B}$. Recall $\mathbf{R}_{\Delta\mathbf{p}^{\prime}}$ in (III-C), $\mathbf{R}_{\Delta\mathbf{p}}$ in (III-B) and $C_{A|B,G}=\log\frac{N_{A|B}}{D_{A|B}}$ in (III-D). Also recall $\mathbf{H}_{AB}^{\prime\prime}=\sqrt{p_{B}/(2n_{B})}\mathbf{H}_{AB}$, $\mathbf{H}_{EB}^{\prime\prime}=\sqrt{p_{B}/(2n_{B})}\mathbf{H}_{EB}$, $\mathbf{H}_{BA}^{\prime}=\sqrt{p_{A}/n_{A}}\mathbf{H}_{BA}$, and $\mathbf{H}_{EA}^{\prime}=\sqrt{p_{A}/n_{A}}\mathbf{H}_{EA}$.
@@
$\lim_{p_{A}\to\infty}\log D_{A|B}=\log\left|\frac{\eta_{p}n_{A}}{2n_{B}}\mathbf{H}_{AB}^{H}\mathbf{H}_{AB}(\boldsymbol{\Pi}_{BA}^{2})^{-1}+\mathbf{I}_{n_{B}}\right|$ (129)
@@
$\lim_{p_{A}\to\infty}\log N_{A|B}=\lim_{p_{A}\to\infty}\log\left|\frac{p_{B}}{2n_{B}}\mathbf{H}_{AB}^{H}\mathbf{H}_{AB}\right|+o(\log p_{A}).$ (130)
@@
Now let us consider $C_{E|B,G}=\log\frac{N_{E|B}}{D_{E|B}}$ in (III-E) and $\mathbf{T}$ in (III-E). We know that $rank(\mathbf{H}_{EA})=\min(n_{A},n_{E})\doteq r_{A}$. We can write the eigenvalue decomposition (EVD) of $\mathbf{H}_{EA}^{H}\mathbf{H}_{EA}$ as $\mathbf{U}\mathbf{D}^{2}\mathbf{U}^{H}=\mathbf{U}_{A}\mathbf{D}_{A}^{2}\mathbf{U}_{A}^{H}$ where $\mathbf{D}_{A}^{2}$ is $r_{A}\times r_{A}$ nonsingular diagonal and $\mathbf{U}_{A}$ is the corresponding $r_{A}$ columns of the $n_{A}\times n_{A}$ unitary matrix $\mathbf{U}$. It follows from (III-E) that
@@
$C_{key}=\log\left|\mathbf{I}_{n_{A}}+\frac{p_{A}}{n_{A}}\mathbf{\tilde{H}}_{EA}^{H}\mathbf{\tilde{H}}_{EA}\right|-\log\left|\frac{p_{A}}{n_{A}}\mathbf{H}_{EA}^{H}\mathbf{H}_{EA}+\mathbf{I}_{n_{A}}\right|$ (145)
@@
Next we consider $R_{s,G}^{B,A}\doteq\lim_{p_{A}\to\infty}R_{s,G}^{B}$. Since $\mathbf{H}_{EA}^{H}\mathbf{H}_{EA}$ is invertible, then $\lim_{p_{A}\to\infty}\mathbf{T}=\mathbf{I}_{n_{B}}$ and $\lim_{p_{A}\to\infty}\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}=0$. Also note that $\mathbf{R}_{\mathbf{\hat{p}}}=\mathbf{I}_{n_{B}}+\mathcal{O}(1/p_{A})$ and $\lim_{p_{A}\to\infty}\mathbf{R}_{\Delta\mathbf{p}^{\prime}}=\lim_{p_{A}\to\infty}\mathbf{R}_{\Delta\mathbf{p}}=0$. Therefore,
$R_{s,G}^{B,A}\doteq\lim_{p_{A}\to\infty}R_{s,G}^{B}=\lim_{p_{A}\to\infty}\log\left(\frac{|\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}|}{|\mathbf{R}_{\Delta\mathbf{p}}|}\right)$ (149)
@@
$R_{s,G}^{B,A}=\lim_{p_{A}\to\infty}\log|\mathbf{R}_{\Delta\mathbf{p}}^{-1}\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}|$
@@
It is easy to verify that subject to (115), $R_{s,1}>0$ if and only if
@@
We will also show explicitly that $t_{1,M}<M-1$. This means that if $S_{E,A}=\alpha_{1}S_{1}>M-2$, then $01-\frac{t_{1,M}}{\alpha_{1}S_{1}+1}>$ 0 and hence (155) is equivalent to (116). Since $R_{s,1}$ must be a decreasing function of $\alpha_{1}$ (which is the ratio of the channel strength from AP to Eve over that from AP to UE1), $R_{s,1}$ must increase as $\alpha_{1}$ decreases. If the peak value of $\frac{t_{1,M}}{\alpha_{1}S_{1}+1}$ (a decreasing function of $\alpha_{1}$) is larger than one, then as $\alpha_{1}$ decreases (starting from the condition $S_{E,A}=\alpha_{1}S_{1}>M-2$) the condition $01-\frac{t_{1,M}}{\alpha_{1}S_{1}+1}>$ 0 would be reversed and a contradiction from $R_{s,1}>0$ would be concluded. Therefore, $t_{1,M}<\alpha_{1}S_{1}+1$ must hold for all $\alpha_{1}$. This also means that $t_{1,M}<1$ when $S_{E,A}=\alpha_{1}S_{1}=0$.
@@
Also recall $c_{i}=\sigma_{\hat{p}_{i}}^{2}=\mathbb{E}\{|\hat{p}_{i}|^{2}\}$; and for $n_{A}=1$, $\phi_{i,j}=1$ and hence $\epsilon_{i,j}=c_{i}c_{j}$.
@@
$q_{b}=-\mathbf{c}_{a}^{H}\mathbf{B}_{1,1}\mathbf{c}_{a}\mathbf{g}_{A}^{H}(\mathbf{g}_{A}\mathbf{g}_{A}^{H}+\mathbf{I})^{-1}\mathbf{g}_{A}=-q_{a}\frac{S_{E,A}}{S_{E,A}+1}.$ (176)
@@
$\gamma_{1}-1=c_{1}-q_{1}=\left(c_{1}-c_{1}^{2}\frac{S_{E,A}}{S_{E,A}+1}\right)-\frac{c_{1}^{2}t_{1,M}}{(S_{E,A}+1)^{2}}$ (201)
@@
-D Proof of proposition 7
Let $\mu=\frac{\sigma^{2}}{1+\sigma^{2}}$, $\mu^{\prime}=1-\mu$, $\mu_{E,A}=\frac{\sigma_{E,A}^{2}}{1+\sigma_{E,A}^{2}}$, $\mu_{E,A}^{\prime}=1-\mu_{E,A}$. It follows from (123) that
$\frac{1}{\sigma_{\Delta s_{i}}^{2}}=1+\frac{1}{g_{A}}$ (202)
with $g_{A}=\mu\mu^{\prime}+\sigma_{A}^{2}$.
Note that $\sigma_{\Delta s_{i}}^{2}$ is invariant to $i$. Also due to symmetry of the network,
$\sigma_{\Delta s_{1|E}}^{2}>\sigma_{\Delta s_{2|E,1}}^{2}>\cdots>\sigma_{\Delta s_{M|E,1:M-1}}^{2}.$ (203)
So, the descending order of the terms in (128) is clear.
We now need to prove $\sigma_{\Delta s_{M|E,1:M-1}}^{2}>\sigma_{\Delta s_{M}}^{2}$ subject to a sufficient power $p_{B}$, we first write
$\mathbf{A}_{M}-\frac{1}{b}\mathbf{c}\mathbf{c}^{T}=\left[\begin{array}[]{cc}\mathbf{B}_{M}&\mathbf{d}\\ \mathbf{d}^{T}&a_{E}\end{array}\right]$ (206)
where $a_{E}=1+\mu^{\prime}+\sigma_{E}^{2}-\mu_{E,A}^{\prime 2}\mu^{\prime 2}$, $\mathbf{d}=\mu_{E,A}\mu^{\prime 2}\mathbf{1}$ and
$\mathbf{B}_{M}=\mu_{E,A}\mu^{\prime 2}\mathbf{1}\mathbf{1}^{T}+(a_{E}-1-\mu_{E,A}\mu^{\prime 2})\mathbf{I}.$ (207)
Applying $(\mathbf{I}+\mathbf{x}\mathbf{x}^{H})^{-1}=\mathbf{I}-\frac{1}{1+\|\mathbf{x}\|^{2}}\mathbf{x}\mathbf{x}^{H}$ to (126) with $i=M$, one can verify that
$\frac{1}{\sigma_{\Delta s_{M|E,1:M-1}}^{2}}=1+\frac{1}{g_{E}}$ (208)
with
$g_{E}=\sigma_{E}^{2}+\mu^{\prime}-\mu^{\prime}_{E,A}\mu^{\prime 2}-\frac{(M-1)\mu_{E,A}^{2}\mu^{\prime 3}}{\frac{a_{E}^{2}}{\mu^{\prime}}+\mu+(M-1)\mu_{E,A}\mu^{\prime}}$ (209)
To compare (202) and (208), we consider
$g_{E}-g_{A}=\sigma_{E}^{2}-\sigma_{A}^{2}+\mu^{\prime}\frac{\mu_{E,A}\sigma_{E}^{2}+\mu_{E,A}\mu^{\prime}\mu}{\frac{\sigma_{E}^{2}}{\mu^{\prime}}+\mu+(M-1)\mu_{E,A}\mu^{\prime}}.$ (210)
It is obvious that if $\sigma_{E}^{2}\geq\sigma_{A}^{2}$, we have $g_{E}-g_{A}>0$ and hence $R_{s,M}^{\prime}>0$ for any (positive) power $p_{B}$ used by each UE.
Now consider the case of $\sigma_{E}^{2}<\sigma_{A}^{2}$ or $\beta_{0}=\frac{\sigma_{A}^{2}}{\sigma_{E}^{2}}>1$. Then $g_{E}-g_{A}>0$ if and only if
$c_{2}\sigma_{A}^{4}+c_{1}\sigma_{A}^{2}-c_{0}<0$ (211)
or equivalently,
$\sigma_{A}^{2}<\frac{1}{2}\left(-\frac{c_{1}}{c_{2}}+\sqrt{\left(\frac{c_{1}}{c_{2}}\right)^{2}+4\frac{c_{0}}{c_{2}}}\right)\doteq\bar{\sigma}_{A}^{2}$ (212)
where $c_{2}=\frac{\beta_{0}-1}{\beta_{0}^{2}\mu^{\prime 2}}$, $c_{1}=\frac{(\beta_{0}-1)(\mu+(M-1)\mu_{E,A}\mu^{\prime})}{\beta_{0}\mu^{\prime}}-\frac{\mu_{E,A}}{\beta_{0}}$, and $c_{0}=\mu_{E,A}\mu\mu^{\prime}$. Here $\bar{\sigma}_{A}^{2}$ is positive and invariant to $\sigma_{A}^{2}$ while $\sigma_{A}^{2}$ is inversely proportional to the power $p_{B}$ used by each UE in phase 2. So, we have proven that subject to any given $\beta_{0}>1$, there is a power threshold $\bar{p}_{B}$ such that the smallest term $R_{s,M}^{\prime}$ in (128) is positive when $p_{B}>\bar{p}_{B}$.
For $\beta_{0}>1$ and a large $M$, we have $c_{1}=\mathcal{O}(M)$ and hence
$\bar{\sigma}_{A}^{2}\approx\frac{1}{2}\left(-\frac{c_{1}}{c_{2}}+\frac{c_{1}}{c_{2}}\left(1+\frac{1}{2}\frac{4c_{0}/c_{2}}{c_{1}^{2}/c_{2}^{2}}\right)\right)=\frac{c_{0}}{c_{1}}$
$=\mathcal{O}\left(1/M\right).$ (213)
In this case, $\bar{p}_{B}$ increases linearly with $M$.